Does the origin lie inside or outside the circle X ^ 2 + Y ^ 2 - 4 x - y=0?
Grade:

1 Answers

Rathod Shankar AskiitiansExpert-IITB
12 years ago

Hi Burhan,

say F(x,y) =

X ^ 2 + Y ^ 2 - 4 x - y

first step is to insert x= 0 and y=0 in this equations if that value comes negative than origin lie inside or outside

F(x,y) = 0+0-0*4-0 =0 

 

if the value is o then origin lies on the circle not inside or outside but on the circle

 

Thank You